### v2

# Investigating Energy Efficiency and Performance Trade-offs in LLM Inference Across Tasks and DVFS Settings

Abstract
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks, leading to widespread adoption in both research and industry. However, their inference workloads are computationally and energy intensive, raising concerns about sustainability and environmental impact. As LLMs continue to scale, it becomes essential to identify and optimize the factors that influence their runtime efficiency without compromising performance. In this work, we systematically investigate the energy-performance trade-offs of LLMs during inference. We benchmark models of varying sizes and architectures, including Falcon-7B, Mistral-7B-v0.1, LLaMA-3.2-1B, LLaMA-3.2-3B, and GPT-Neo-2.7B, across tasks such as question answering, commonsense reasoning, and factual generation. We analyze the effect of input characteristics, such as sequence length, entropy, named entity density and so on. Furthermore, we examine the impact of hardware-level optimizations through Dynamic Voltage and Frequency Scaling (DVFS), measuring how different GPU clock settings affect latency and power consumption. Our empirical findings show that model architecture, input complexity, and clock configuration significantly influence inference efficiency. By correlating input features with energy metrics and evaluating DVFS behavior, we identify practical strategies that reduce energy consumption by up to 30% while preserving model quality. This study provides actionable insights for designing energy-efficient and sustainable LLM inference systems.
I Introduction
The large-scale pre-trained models like GPT, BERT, and T5 are increasingly deployed in real-world scenarios for natural language processing (NLP) applications such as text generation, question answering, and machine translation, among others. The rapid growth in the adoption of such large scale models has created immense demand for computing resources and energy systems. Therefore, such models should be optimized for performance and energy consumption for sustainable AI development. Accordingly, it is crucial to understand the performance and energy efficiency of these models across diverse tasks and input characteristics, and hardware settings.
Recent trends in the development of large language models (LLMs) have primarily focused on increasing the number of parameters to enhance model capabilities and improve predictions by incorporating larger training datasets. Early models, such as BERT [1] and GPT-2 [2], had several hundred million parameters. Today’s top-performing models, like LLaMA 3 [3] and GPT-4 [4], consist of hundreds of billions of parameters. Alongside the increase in size, domain-specific fine-tuned smaller versions of models (i.e., small LLMs) are also becoming increasingly capable.
Recent advancements in NLP have focused primarily on improving the accuracy and generalization capabilities of these models. However, as the models scale up in size and complexity, their computational demands have risen dramatically, leading to increased energy consumption during both training and inference [5, 6, 7]. For instance, GPT-3, with 175 billion parameters, consumed an estimated 1,287 MWh of energy during training [8], and its inference processes remain energy-intensive, particularly when handling large-scale tasks. While training energy costs are one-time, inference is a continuous process as models are deployed to serve millions of queries daily. For example, GPT-3 can consume 0.0003 kWh per query during inference, and when scaled to millions of users, this energy consumption becomes a significant challenge. As a result, there is a growing need to explore energy-performance trade-offs, especially during inference.
Existing research has explored optimizations in model architectures, such as model pruning, quantization, and distillation, to reduce computational costs. For instance, model distillation has been shown to reduce model sizes by up to 90%, leading to a 50-60% reduction in energy consumption during inference [9]. Similarly, techniques like model pruning and quantization can improve inference efficiency by reducing precision and eliminating unnecessary model weights, leading to faster inference times and lower power consumption [10]. However, much of the existing work has focused on general optimizations rather than how specific NLP tasks or input characteristics affect model performance and energy efficiency.
To address this, we seek to answer two critical research questions. First, we analyze how different types of NLP tasks, such as text generation, question-answering, and logical reasoning tasks, impact the energy consumption and performance of large pre-trained models during inference. Also, we investigate how input data characteristics like sequence length, token entropy, entity count so on affect energy consumption. Second, we explore the impact of hardware-based power-saving techniques, such as Power Capping and Dynamic Voltage Frequency Scaling (DVFS), on model’s energy efficiency and performance. By systematically examining these factors, we aim to provide insights into optimizing LLMs for specific tasks and improving the overall sustainability of NLP systems.
In this study, we present a comprehensive evaluation of large language models (LLMs), analyzing their inference-time performance and energy consumption across diverse architectures, task types, and hardware settings. We benchmark five state-of-the-art models ranging from 1 billion to 7 billion parameters on tasks such as binary question answering, commonsense reasoning, and open-ended generation using an NVIDIA A100 GPU. Our analysis highlights that larger models such as Mistral 7B and Falcon 7B achieve higher accuracy, particularly on complex tasks, but incur up to six times more energy and memory usage compared to efficient models like LLaMA 3.2 1B. To uncover the sources of computational variation, we examine eight input-level features including sequence length, entropy, and named entity density. We show that certain attributes, such as longer inputs or higher lexical diversity, directly correlate with increased energy cost. Furthermore, by leveraging dynamic voltage and frequency scaling (DVFS), we show that tuning the GPU’s SM clock, can significantly reduce inference time and improve energy efficiency by up to 30% without requiring any modifications to the model.
These findings emphasize the importance of aligning model architecture, task requirements, and hardware configurations to achieve optimal trade-offs between performance and energy efficiency. By focusing on task-specific optimization and leveraging hardware capabilities effectively, it is possible to enhance the sustainability of large language models while maintaining their utility across diverse NLP applications.
II Methodology
To evaluate energy-performance trade-offs in large language models (LLMs), we design a benchmarking framework covering diverse architectures and NLP tasks. Our methodology captures energy usage, performance, and input-level complexity under controlled inference settings. The following sections describe the selected models, datasets, and experimental setup.
| Model | Params | Type | Primary Use Cases |
| tiiuae/falcon-7b [11] | 7B | Decoder (GPT) | Text generation, extended context handling |
| Mistral-7B [12] | 7B | Decoder | Long-context generation, multilingual tasks |
| EleutherAI/gpt-neo-2.7B [13] | 2.7B | Decoder (GPT) | General-purpose language tasks, reasoning |
| meta-llama/LLaMA-3.2-3B [14] | 3.2B | Decoder | Low-latency inference, instruction tuning |
| meta-llama/LLaMA-3.2-1B [15] | 1.2B | Decoder | Lightweight deployment, efficient QA |
II-A Model and Dataset Selection
We benchmark five pre-trained decoder models: Falcon-7B, Mistral-7B, LLaMA-3.2-1B, LLaMA-3.2-3B, and GPT-Neo-2.7B, selected for their diversity in size (1B–7B parameters) and architecture (Table I). These models balance recent advancements and deployment efficiency, enabling a representative comparison of energy and performance trade-offs. Evaluation is performed across tasks such as classification, question answering, and commonsense reasoning, summarized in Table II, to ensure coverage of both simple and complex inference scenarios.
| Dataset | Task Type | Evaluation Metric |
| BoolQ [16] | Binary QA | Accuracy |
| HellaSwag [17] | Commonsense MCQ | Accuracy |
| Winogrande [18] | Pronoun Resolution (MCQ) | Accuracy |
| TruthfulQA (Gen) [19] | Open-ended QA (Generation) | ROUGE-1 [20] |
II-B Testbed Setup
All experiments were conducted on an NVIDIA A100 80GB PCIe GPU with 512GB of RAM, running Ubuntu 22.04. The A100 features 80 GB of high-bandwidth memory (HBM2e), a 300W TDP, and supports extensive configurability through Dynamic Voltage and Frequency Scaling (DVFS). Specifically, we varied the Streaming Multiprocessor (SM) clock frequency between 210 MHz and 1410 MHz, and the memory clock was fixed at 1215 MHz throughout the experiments. This allowed fine-grained control over performance and power consumption, in line with prior DVFS studies [21]. Energy usage was measured using nvidia-smi via periodic polling during inference runs. All models were implemented using PyTorch and executed using the Hugging Face Transformers library to ensure reproducibility and consistency across benchmarks.
II-C Evaluation Metrics
We evaluate task performance using Accuracy for classification and binary QA tasks (BoolQ, HellaSwag, Winogrande), and ROUGE-1 for generation (TruthfulQA). Energy consumption (Joules) is measured as the product of average power and total inference time. Throughput is the number of tokens processed per second, indicating the model’s inference speed.
To analyze input-level complexity, we compute eight features that capture linguistic and semantic properties. Sequence length reflects the total number of tokens, serving as a proxy for processing workload. Entropy [22] quantifies token diversity and unpredictability. Readability score (Flesch-Kincaid grade [23]) measures syntactic complexity and ease of comprehension. Named entity count indicates factual density by measuring the number of real-world entities. Embedding variance [24] captures representational diversity, while self similarity [25] reflects textual redundancy via average cosine similarity between sentence pairs. Contextual coherence [26] quantifies narrative flow based on the similarity between the first and last sentence embeddings. Finally, model perplexity [27] serves as a proxy for model-level difficulty. We correlate each feature with per-query energy to examine their influence on inference cost.
III Benchmarking and Characterization of LLM inference
Understanding how LLMs behave across different types of NLP tasks is essential for designing efficient inference pipelines. This section presents a detailed benchmarking study that captures the interplay between task type, model behavior, and system-level performance.
III-A Characterizing LLM’s Task Diversity
The benchmarking of large language models (LLMs) across diverse tasks provides critical insights into their performance, energy efficiency, and resource utilization. Tasks such as Boolean classification, question answering and commonsense reasoning are evaluated to highlight the trade-offs between model accuracy, latency, energy consumption, and memory usage. The study highlights how task complexity and model architecture impact efficiency, stressing the need to align model choice with application goals. All tasks were run with a batch size of 8 across 600 queries, repeated three times for averaging.
III-A1 Energy Consumption
Mistral-7B consistently exhibits the highest energy consumption across all tasks, ranging from approximately 2,800 to 3,100 joules per 1,200 queries (Fig 1). In contrast, LLaMA-3.2-1B operates with remarkable energy efficiency, consuming only about 650–730 joules for the same query load, a 77% reduction compared to Mistral-7B. Falcon-7B also demonstrates better energy efficiency than Mistral, averaging around 2,650–2,900 joules. GPT-Neo-2.7B is notably inefficient; despite having fewer parameters than Falcon and Mistral, it consumes up to 6,968 joules on BoolQ which is more than double the energy used by models with higher performance. This highlights that architectural inefficiencies of such larger unoptimized KV caches and slower attention mechanisms that can dominate over parameter count in dictating energy cost.
III-A2 Performance Metric
In classification tasks such as BoolQ and Winogrande, Mistral-7B achieves the best accuracy, scoring 81.3% and 73.8%, respectively (Fig 2). Falcon-7B trails closely, with accuracies between 69% and 72% across tasks. GPT-Neo-2.7B underperforms across the board, achieving just 62.7% on BoolQ and 57.1% on Winogrande. LLaMA-3.2-1B, while being the most efficient, pays a substantial performance penalty up to a 27 percentage point drop in accuracy on BoolQ compared to Mistral-7B. On the generative TruthfulQA-Gen task, Mistral-7B again leads with a ROUGE-1 score of 38.4%, outperforming all other models by a margin of over 6%.
III-A3 Peak Memory Usage
Peak memory usage mirrors model size and attention design. Falcon-7B and Mistral-7B consistently require 14.7–14.8 GB across all tasks (Fig 3). LLaMA-3.2-3B consumes approximately 6.9–7.2 GB, offering a 50% reduction. LLaMA-3.2-1B is the most memory-efficient, using just 3.2 GB, an impressive 78% lower than Mistral-7B. In contrast, GPT-Neo-2.7B exhibits relatively high memory usage (ranging from 10.9 to 11.3 GB) despite delivering lower performance, highlighting an inefficient trade-off between energy consumption, memory usage, and accuracy.
III-A4 Throughput
In terms of tokens processed per second, LLaMA-3.2-1B is a clear outlier in performance (Fig 4). It delivers over 30,000 tokens per second on BoolQ which is roughly 3.3 times faster than Mistral-7B and 6 times faster than GPT-Neo-2.7B. Even on more complex generative tasks like TruthfulQA-Gen, it maintains a throughput above 1,000 tokens/sec, highlighting its scalability for high-throughput use cases. Mistral-7B and Falcon-7B average around 10,000 tokens/sec on classification tasks and drop to about 500–600 tokens/sec on generative workloads, indicating increased KV cache and attention overhead with longer sequences.
III-B Impact of Input Characteristics on Inference Efficiency
Understanding how input-level complexity affects LLM inference is essential for optimizing energy usage and latency. We investigate this by analyzing eight linguistic and semantic features across two contrasting tasks: BoolQ and HellaSwag. Each input is grouped into feature bins, and we compute the average energy consumed per query to assess how these features influence computational cost.
III-B1 Input Feature Complexity and Energy Implications
To uncover how input level characteristics influence energy consumption and model behavior during LLM inference, We organize the input features into three categories:
Textual Properties: word_length, entropy, readability, and named_entities
Embedding-Based Metrics: embedding_variance, self_similarity, and contextual_coherence
Model Difficulty Estimate: perplexity
Our correlation analysis (figures 7 and 8)) reveals several consistent patterns. The strongest relationship observed is between word_length and entropy ( in HellaSwag, in BoolQ) suggesting that these metrics encode overlapping aspects of input complexity. Including both in regression or classification pipelines may introduce multicollinearity. In HellaSwag, readability correlates moderately with length () and entropy (), indicating that longer and lexically richer texts tend to have more complex syntactic structure. In BoolQ, however, readability exhibits a weak correlation with length () and entropy (), emphasizing its limited role in modeling QA-style inputs. Named_entities shows a task-specific pattern. It correlates strongly with length in BoolQ (), where longer passages usually contain more entities, consistent with the factual nature of QA content. In HellaSwag, the same correlation is lower (), as narrative texts may emphasize coherence over explicit factual density.
Among embedding-derived features, embedding_variance and self_similarity are perfectly inversely correlated (), and both show strong relationships with contextual_coherence (variance , self-similarity ). These patterns indicate that only one embedding metric is necessary to describe sentence-level representation variability.
Perplexity, which is used as a proxy for model confidence, negatively correlates with input length and entropy in both datasets. In HellaSwag, the correlation is strong ( with length, with entropy), while in BoolQ the values are more modest ( with length, with entropy). The observed patterns suggest that inputs lacking sufficient length or semantic detail result in greater prediction uncertainty, reflected in higher perplexity.
Our energy-based analysis confirms these trends (figures 6 and 5). For example, in HellaSwag, queries in the highest length bin (bin 4: 88.0–116.0 tokens) consumed an average of 10.25 J/query on GPT-Neo-2.7B, while the lowest bin (11.0–36.0 tokens) required just 5.12 J/query, that is over 2× energy difference. A similar upward trend is observed for named_entities and entropy, with GPT-Neo-2.7B and Falcon-7B also showing increasing energy with input complexity. Interestingly, bins with higher perplexity (e.g., bin 0: 4.68–20.14) showed lower average energy usage (~3.1 J/query), supporting the idea that models expend less effort when uncertainty is high and response length is short.
Models exhibit different sensitivities to input characteristics. Falcon-7B shows consistent energy scaling across length and entropy bins, whereas LLaMA-3.2-1B remains flat, with ~1.2-1.4 J/query across all bins, making it suitable for lightweight inference. Embedding-based features, on the other hand, show limited impact on energy usage. The variance across bins was < 0.3 J/query for all models, therefore reinforcing their marginal utility in energy-aware systems.
These findings suggest that input features not only determine linguistic and semantic complexity but also significantly influence computational behavior. Selecting the right features is key to designing inference-aware architectures and adaptive execution strategies.
III-C Dynamic Voltage Frequency Scaling (DVFS)
Modern GPUs like the NVIDIA A100 support dynamic voltage and frequency scaling (DVFS), enabling fine-grained control over power and performance through configurable clock settings [28]. This feature is particularly crucial for large language models (LLMs), which exhibit varied performance and energy behaviors across tasks [29]. In our study, we selected seven representative SM clock frequencies, 210 to 1410 MHz, to explore the trade-offs between energy and latency. As shown in Figure 9, different models respond differently to DVFS settings, highlighting the importance of task and model-aware tuning. All experiments used a batch size of 8, evaluated over 300 queries per clock level, averaged over three runs.
On the BoolQ task, all models demonstrated substantial latency reductions ranging from 70% to over 80%, indicating strong responsiveness to clock frequency. Among them, GPT-Neo-2.7B showed the most pronounced latency improvement () alongside the steepest increase in power (), yielding a power slope of W/MHz, which is the highest across all tasks and models. In contrast, LLaMA-3.2-1B exhibited the lowest sensitivity, with a modest rise in power and a relatively flat power slope ( W/MHz), making it the most energy-efficient model under frequency scaling. Interestingly, while both Falcon-7B and Mistral-7B achieved similar latency gains ( to ), Mistral incurred a notably steeper power increase, suggesting less efficient voltage scaling.
In the HellaSwag task, a similar trend emerged. GPT-Neo-2.7B again showed the steepest Time/Clock ( s/MHz) and highest power gain (), confirming its aggressive scaling behavior. Meanwhile, LLaMA-3.2-1B maintained its position as the most efficient model, with a relatively shallow power slope and a moderate latency improvement of 62%. Both Falcon and Mistral showed moderate scaling behavior, but Mistral again demonstrated slightly steeper power sensitivity.
The Winogrande dataset revealed a narrower spread in scaling metrics, with latency improvements between 60% and 78% and power increases under 170% for all models. LLaMA models remained the most power-efficient, with Power/Clock values of and W/MHz for the 1B and 3B variants, respectively. GPT-Neo continued to show high power growth, though the overall energy cost per latency gain was lower than in BoolQ or HellaSwag, reflecting the task’s smaller input lengths and lighter reasoning demands.
For TruthfulQA, a generative task with long outputs, the absolute inference time savings were the highest, dropping from over 290 seconds to under 100 seconds for some models. Mistral and GPT-Neo once again showed steep time slopes ( and s/MHz, respectively), but at a high energy cost. Mistral’s power rose by , while GPT-Neo’s power more than doubled (). LLaMA-3.2-1B exhibited both low power and time slopes, achieving a reduction in latency with only a power increase, reinforcing its suitability for energy-constrained generative tasks.
| Task | Model | Clock (MHz) | Power (W) | Time (s) | Energy (J) | Performance metric |
| BoolQ | Falcon-7B | 1005 | 131.47 | 12.05 | 1584.36 | 0.7267 |
| Mistral-7B | 1005 | 130.80 | 12.11 | 1584.55 | 0.8467 | |
| LLaMA-3.2-1B | 1215 | 86.50 | 3.13 | 270.50 | 0.6533 | |
| LLaMA-3.2-3B | 1005 | 97.70 | 6.22 | 608.01 | 0.7733 | |
| GPT-Neo-2.7B | 1005 | 148.45 | 28.63 | 4250.53 | 0.6433 | |
| HellaSwag | Falcon-7B | 1005 | 119.17 | 29.48 | 3513.53 | 0.5200 |
| Mistral-7B | 1005 | 119.53 | 28.28 | 3380.36 | 0.5400 | |
| LLaMA-3.2-1B | 810 | 79.82 | 10.80 | 861.77 | 0.4600 | |
| LLaMA-3.2-3B | 1005 | 94.33 | 17.92 | 1690.78 | 0.4933 | |
| GPT-Neo-2.7B | 1005 | 139.86 | 53.94 | 7544.46 | 0.4600 | |
| Winogrande | Falcon-7B | 810 | 100.00 | 15.73 | 1572.79 | 0.7067 |
| Mistral-7B | 1005 | 107.53 | 14.34 | 1542.24 | 0.7500 | |
| LLaMA-3.2-1B | 810 | 65.40 | 5.55 | 362.75 | 0.5633 | |
| LLaMA-3.2-3B | 1005 | 88.92 | 8.87 | 788.95 | 0.7000 | |
| GPT-Neo-2.7B | 1005 | 120.27 | 17.42 | 2094.77 | 0.5700 | |
| TruthfulQA | Falcon-7B | 1005 | 94.84 | 109.95 | 10427.45 | 0.2667 |
| Mistral-7B | 1005 | 98.14 | 113.74 | 11162.64 | 0.4367 | |
| LLaMA-3.2-1B | 810 | 63.55 | 47.10 | 2993.50 | 0.2067 | |
| LLaMA-3.2-3B | 1005 | 76.45 | 61.68 | 4715.46 | 0.1933 | |
| GPT-Neo-2.7B | 1005 | 102.84 | 103.08 | 10600.86 | 0.3733 |
III-C1 DVFS-Aware Efficiency Trends
As shown in table III, across all evaluated tasks, LLaMA-3.2-1B consistently demonstrated the lowest energy consumption, achieving values as low as 270.5 J on BoolQ, 861.8 J on HellaSwag, and 362.8 J on Winogrande. This efficiency is primarily attributed to its compact size and effective DVFS scaling at mid-range SM clock frequencies (810–1215 MHz). However, this comes with a trade-off in terms of moderate performance, with accuracy ranging from 0.6533 on BoolQ to 0.2067 on TruthfulQA, suggesting that LLaMA-3.2-1B is best suited for edge deployments or energy-constrained environments where reduced accuracy is acceptable.
In contrast, Mistral-7B emerged as the best energy-accuracy balanced model. It achieved the highest task-specific accuracy: 0.8467 on BoolQ, 0.7500 on Winogrande, and 0.4367 on TruthfulQA, while maintaining moderate energy usage (1542.2 J, 1542.2 J, and 11162.6 J, respectively). These results show that Mistral-7B offers a strong middle ground for use cases demanding both performance and efficiency, making it ideal for power-aware, real-time inference applications.
On the other end of the spectrum, GPT-Neo-2.7B consistently incurred the highest energy consumption across tasks, with values reaching 4250.5 J on BoolQ, 7544.5 J on HellaSwag, and 10600.9 J on TruthfulQA, while delivering only modest accuracy gains. This indicates that GPT-Neo’s aggressive scaling comes with poor energy proportionality, making it less suitable for constrained deployments where energy efficiency is a key concern.
The results highlight that task characteristics strongly impact DVFS efficiency. Generative tasks like TruthfulQA incur high energy costs even at optimal clocks due to longer runtimes, while classification tasks (e.g., BoolQ, Winogrande) benefit from mid-range frequencies (810–1005 MHz). This emphasizes the need for task-aware clock tuning over static DVFS settings.
III-C2 Contrasting DVFS Trends in LLMs vs. Conventional Workloads
Prior work on GPU DVFS shows that energy savings depend heavily on workload characteristics. Mei et al.[30] and Ge et al.[31] found that lowering core and memory frequencies often reduces energy use in traditional GPU workloads like matrix multiplication, FFT, and memory-bound tasks. Similarly, Tang et al.[32] noted that DNNs achieve optimal efficiency at mid-range frequencies, balancing performance and power. However, these findings do not generalize to LLM inference. We find that higher SM clock frequencies (e.g., 1005MHz) consistently improve energy efficiency because the reduction in runtime outweighs the increase in power consumption. This challenges prior assumptions that lower frequencies are always more energy-efficient. While learning-based DVFS strategies like DVFO (Zhang et al. [33]) and hybrid solutions such as EdgeBERT [34] use early exits or control agents, we show that simple manual SM clock tuning alone can yield substantial energy gains in LLM inference without altering model architecture. We further find that DVFS sensitivity varies even across similarly sized models, emphasizing that architectural choices, not just parameter count, drive energy-performance trade-offs.
IV Scope and Threats to validity
This study analyzes the energy and performance dynamics of LLMs across diverse NLP tasks using five models on an NVIDIA A100 GPU. While the insights are broadly applicable, results may vary with different hardware, model architectures, or training data. Key threats to validity include hardware dependence, model-specific optimizations, and variability in pretraining corpora. Metrics like ROUGE and accuracy may not fully capture task-specific performance, and differences in batch size or tokenization could affect results. Moreover, only static DVFS configurations and a limited set of tasks are evaluated, suggesting the need for broader, dynamic, and hardware-aware analysis in future work.
V Related work
Early works on LLM benchmarking emphasized the importance of evaluating task diversity to comprehensively assess model performance across tasks, such as question answering, summarization, and commonsense reasoning[35, 36]. Prior studies primarily focused on performance metrics, including latency, throughput, and accuracy[37, 38]. However, these works often analyzed generic tasks without extensive consideration of challenges posed by LLM inference, such as energy efficiency and memory utilization.
Task-specific performance characterizations have been explored in works focusing on tasks like summarization and question answering, highlighting that larger models often excel in performance metrics at the cost of increased energy consumption and latency [39, 6, 40]. However, prior studies do not systematically analyze the trade-offs between architectural choices and task requirements.
Dynamic voltage and frequency scaling (DVFS) has gained traction in optimizing energy efficiency for LLM inference [41]. It has been observed that mid-to-high clock configurations achieve optimal energy efficiency by balancing runtime and power consumption, particularly in tasks requiring diverse computational demands[42]. Yet, many of these analyses focus on generic tasks or omit the joint characterization of task-specific energy consumption and model performance [43].
In contrast to these studies, our work provides a holistic evaluation of LLM inference by jointly analyzing task diversity, energy consumption, and hardware utilization. Unlike earlier works that typically focus on a single type of task or model, we explore a broad spectrum of tasks from Boolean classification to commonsense reasoning and compare the performance of models ranging from lightweight architectures like LLaMA-3.2-1B to mid-sized models like Falcon-7B. Additionally, our work extends prior analyses by incorporating input sequence length intervals, revealing nuanced trade-offs between model performance and resource efficiency.
VI Conclusions
This study highlights that LLM inference efficiency is shaped by both model architecture and input characteristics. Larger models like Mistral-7B achieve higher accuracy but consume up to 6× more energy than compact models like LLaMA-3.2-1B. Notably, our results show that setting the GPU SM clock to 1005 MHz often yields the best energy-to-performance trade-off, enabling energy savings of up to 30% compared to the default clock. Input features such as sequence length and entropy strongly correlate with energy use, but due to their high mutual correlation, using just one suffices. Named entity count and perplexity capture orthogonal complexity aspects, while embedding-based features offer limited practical value in energy-aware inference. Overall, careful model selection, manual GPU clock tuning, and lightweight input-aware strategies offer significant opportunities for sustainable and efficient LLM deployment across diverse NLP tasks.